How many seconds will light leaving Los Angeles take to reach the following locations (a) San Francisco (about 500km), (b) London (~10,000km), (c) the Moon (400,000km), (d) Venus (0.3 A.U. from Earth at its closest approach).
Answer by Alan3354(69443)

How many seconds will light leaving Los Angeles take to reach the following locations (a) San Francisco (about 500km), (b) London (~10,000km), (c) the Moon (400,000km), (d) Venus (0.3 A.U. from Earth at its closest approach).
--------------
The commonly used term for the speed of light is 3e5 km/sec.
----
(a) San Francisco (about 500km)
d = r*t
t = d/r = 500/3e5 =~ 1.67e-3 seconds = 0.00167 seconds or 1.67 milliseconds.
=========
Do the others the same way.
===========
PS blinking your eyes takes 100 to 400 milliseconds
In 100 milliseconds light travels 30,000 km or 18,000 miles